# Measuring Casimir Force Across a Superconducting Transition

### v1

@@ IV Experimental Setup and Measurement Protocol
To isolate superconductivity-induced Casimir shifts from material effects and thermal expansion, we use two chips with distinct gaps: a small-gap cavity (190 nm) where Casimir forces are strong, and a large-gap cavity (1213 nm) where they are negligible. Mounting both on the same holder (Fig. 3c) ensures identical environmental conditions, enabling subtraction of material-dependent effects. Although the NbTiN films are co-deposited, we verify that both chips exhibit matched mechanical frequencies and $T_{C}$ (see Supplementary Information A). This differential approach allows us to subtract shared material responses across $T_{C}$ and isolate Casimir forces.
Beyond intrinsic material changes, all other extraneous forces must be minimized or eliminated, as detailed in the Methods section. Electrostatic interactions between plates are canceled via DC backgate biasing, while tip-induced forces are compensated through STM feedback. Van der Waals contributions are suppressed by tuning the tunneling current to stay in a regime where such effects are negligible. Once these known extraneous forces are removed, temperature sweeps around $T_{C}$ isolate the remaining Casimir signal.
Short-Range Measurements Around $T_{C}$
To probe superconductivity-induced variations in the Casimir force, we performed high-precision frequency measurements across $T_{C}\sim 14.2$ K. Each run began at 4.45 K, with controlled heating and stabilization at each step to minimize drift and maintain STM tip stability.
Each sweep across $T_{C}$ lasted up to 16 hours with continuous stabilization and data acquisition. While open-flow helium cooling is resource-intensive (0.45 L/hour), it eliminates mechanical vibrations that would otherwise degrade STM performance and our experiments. The STM holds 9 L of liquid helium — 1.8 L for overnight tip and temperature stabilization, and 7.2 L for the $T_{C}$ sweep. Ultra-high vacuum ($\sim 10^{-10}$ mbar) ensures a clean surface after annealing and suppressed squeezed-film damping at nanoscopic separations.
After systematically eliminated electrostatic forces between the plates and forces between the tip and membrane, we tracked membrane resonance frequency shifts, isolating Casimir contributions by comparing small-gap (190 nm) and big-gap (1213 nm) samples. Big-gap samples (with negligible Casimir force) allow us to account for any unknown changes in frequency coming from material changes at $T_{C}$, thermal expansion, or any other non-Casimir related changes. Fig. 4a shows sharper frequency shifts $\Delta f$ in the small-gap cavity at $T_{C}\sim 14.2$ K, whereas the shifts in the big-gap cavity serve as a reference and capture shifts due to variations in mechanical properties as temperature sweeps. For consistency across runs, we define the reference value as the average frequency near 13.9 K. We use $\Delta\psi$ to denote temperature-dependent variations, and $\delta\psi$ for fixed-temperature changes in quantities $\psi$ such as frequency or pressure. Independent resistivity measurements confirmed consistent $T_{C}\sim 14.2$ K across both small- and big-gap devices (Fig. 4d), using forward and backward sweeps.
Following the calibration in Supplementary Information G, we linearly fit the thermal expansion of the superconductor using data below $T_{C}$ (Fig. 4a), allowing determination of the residual frequency square shift $\Delta\omega^{2}$ above $T_{C}$ for membranes with different gaps, where $\omega=2\pi f$ is the angular resonance frequency. Since the gap size is the only difference between membranes, we attribute the contrasting $\Delta\omega^{2}$ behavior to a separation-dependent force. With other distance-dependent forces minimized, we attribute the frequency shift to changes in the Casimir force. Using $\Delta\omega^{2}$ from the big-gap device, we estimate NbTiN’s mechanical changes and compute the Casimir-induced $\omega^{2}$ shift in the small-gap membrane. The relation $\Delta\omega^{2}=-\Delta P^{\prime}(d)/\rho t$ links frequency shift to the pressure gradient, where $d$ is the separation, $\rho$ the density, and $t$ the thickness (Supplementary Information F). We extract a Casimir pressure gradient variation of $12.10\pm 1.38$ kPa/m for a $0.20\pm 0.02$ K temperature change (Fig. 4c).
In our on-chip parallel-plate setup, the combination of large area and small gap $d$ yields a strong Casimir force, enabling high sensitivity to small variations [37]. This amplifies superconductivity-induced changes, resulting in a pronounced fractional frequency shift that scales with the pressure and force gradients: $\Delta f/f\approx-\Delta P^{\prime}_{\rm Cas}L^{2}/(4\pi^{2}\sigma t)=-\Delta F^{\prime}_{\rm Cas}/(4\pi^{2}\sigma t)$ (Supplementary Information F).
Based on Lifshitz theory [14], we estimate that near $T_{C}$, the Casimir pressure on the suspended NbTiN membrane at a 190 nm separation fits as $P_{\rm Cas}|_{d=190\rm nm}=-1.081\times 10^{-24}/d^{3.507}$ [Pa] $\approx-0.403$ Pa (with $d$ in meters; negative sign denotes attraction). Using this fit, a finite element simulation estimated that the measured $\omega^{2}$ variation (Fig. 4b) corresponds to a Casimir force change of $\Delta F_{\rm Cas}=-0.33\pm 0.04$ nN, or equivalently, a pressure change of $\Delta P_{\rm Cas}=-0.65\pm 0.07$ mPa across the transition. The negative signs indicate that both force and pressure increase as temperature rises through $T_{C}$. Further studies across different gap sizes are needed to clarify this behavior.
Figure 4: (a) Membrane frequency vs. temperature near $T_{C}\approx 14.2$ K for one small-gap sweep and two big-gap sweeps. (b) Frequency square shift $\Delta\omega^{2}$ after calibrating out the temperature dependence of the elastic force (see Supplementary Information G), showing contrasting behavior between small- and big-gap membranes. (c) Pressure gradient variation for the 190 nm-gap membrane. Data (blue dots) are compared with Casimir pressure predictions using the BCS model for ($T<T_{C}$) and either Drude (red) or plasma (purple) models for ($T>T_{C}$) (Supplementary Information H). (d) Forward and backward resistivity sweeps confirm ($T_{C}\sim 14.2$ K) for both membranes.
The abrupt change in Casimir pressure gradient across $T_{C}$ in Fig. 4c was not predicted by existing theory [14]. Since the Casimir force is the derivative of the Helmholtz free energy, $F_{\rm Cas}=-\partial{\cal F}/\partial d$ [38], it is expected to remain continuous through a second-order phase transition. Furthermore, the smooth convergence of BCS and Drude permittivity models at $T_{C}$ supports a continuous force (see Supplementary Information H).
A possible explanation links this behavior to the longstanding Drude–plasma model debate, which centers on the zero-frequency ($n=0$) Matsubara term in the Lifshitz formula. The Drude model predicts a vanishing transverse electric (TE) contribution, while the plasma model allows a finite one that significantly affects the force. While room-temperature experiments favor the plasma model, confirmation in cryogenic regimes — especially for superconductors — remains lacking. To reconcile our observation, we propose a hybrid model using the plasma prescription for $T>T_{C}$ and the BCS permittivity for $T<T_{C}$, resulting in a changing zero-frequency term and a discontinuous pressure gradient (the “Plasma–BCS” model). For comparison, we also consider a Drude–BCS model, which yields a continuous gradient. Both predictions appear in Fig. 4c. Notably, only the Plasma–BCS model reproduces the magnitude of our experimental signal, suggesting it may capture the essential physics.
In Fig. 4c we use $\Omega=5.33$ $\hbar$, consistent with NbTiN estimates [14], but reproducing the full jump requires $\Omega=12.7$ $\hbar$. Because the zero-frequency contribution is poorly understood, $\Omega$ need not match the infrared-derived plasma frequency. A full discussion is provided in Supplementary Information H. If confirmed, this jump would mark a major advance — representing the first measurement sensitive to a transition in the zero-frequency permittivity term across a phase transition, and deepening our understanding of dissipation and quantum fluctuations.
Our results highlight the importance of extreme parallelism, compliant mechanics, advanced material engineering, and high-precision readout in isolating subtle Casimir force variations. While initial findings are promising, further studies are needed with finer temperature sweeps, varying gap sizes, thicknesses, and superconducting materials to better constrain theory. Future work will focus on reproducibility and model refinement, leveraging our platform’s unique sensitivity to probe Casimir–superconductivity interactions.
V Conclusion and Outlook
The interplay between Casimir forces and superconductivity remains an open frontier in quantum physics, with implications for longstanding debates on electron physics, high-$T_{C}$ superconductivity, and perhaps quantum gravity. We introduced a microchip-based platform that achieves extreme parallelism and mechanical stability, enabling high-precision Casimir force measurements across $T_{C}$. Compared to previous superconducting Casimir experiments across $T_{C}$ [16, 17], our system increases the interacting surface area by over three orders of magnitude and achieves pressure resolutions 100 times greater. The generated Casimir forces exceed the average of all previously published Casimir measurements by more than two order of magnitude (Supplementary Information J).
Our approach eliminates thermal, electrostatic, and van der Waals contributions while operating at cryogenic temperatures. We introduce a novel STM readout that directly detects membrane resonance, extending STM’s role from atomic imaging to mesoscopic force sensing. Its highly localized interaction area minimizes disruption to superconductivity and allows precise cancellation of tip-induced forces.
Initial measurements reveal frequency shifts in small-gap membranes near $T_{C}$, consistent with superconductivity-modulated Casimir interactions. If due to Casimir effects, this would be the first evidence that superconductivity alters vacuum fluctuations. The sign of the shift suggests a decrease in Casimir pressure across $T_{C}$, contrary to expectations, and the abruptness appears inconsistent with a second-order phase transition. Existing theory [14] predicts a smooth change, while a model using the plasma prescription for the $n=0$ Matsubara term in the normal state [39, 40, 41, 42, 43] reproduces our result to similar order of magnitude. As this prescription aligns with room-temperature Casimir data, its extension to the superconducting regime is a plausible hypothesis. Nonetheless, broader sweeps and gap variation are needed to verify the trend and rule out alternative causes.
Though our approach sets a new benchmark, further gains in parallelism are limited by stiction from residual charge, motivating collapse recovery methods [44] and the exploration of higher-$T_{C}$ superconductors. Beyond Casimir physics, our platform offers broad utility for sensing, table-top quantum gravity [45], and condensed matter experiments [46]. Its extreme parallelism enhances responsivity via Casimir forces [37], enabling applications in accelerometry [47], optomechanics [48], and other force-sensitive systems.
These results highlight the potential of combining on-chip nanomechanics and STM to probe fundamental quantum effects — opening a new experimental regime where superconductivity and vacuum fluctuations begin to interact.
VI Methods
VI.1 STM measurement protocol
@@
After fabricating the solid-state parallel-plate cavities, we use a scanning tunneling microscope (STM) to measure membrane dynamics with atomic-scale precision. While STM is typically used for atomic imaging and transport studies on ultra-flat surfaces, its extreme sensitivity to tip–sample distance makes it ideal for detecting nanomechanical motion. The tunneling current depends exponentially on separation, with 0.1 nm changes producing an order-of-magnitude variation. Here, we repurpose STM to track the mechanical resonance of suspended superconducting NbTiN membranes, extending its use from atomic-scale measurements to mesoscopic devices.
@@
Two aspects of the STM’s atomic-scale precision are critical for our measurements. First, its high vertical ($z$) resolution enables detection of membrane resonance by tracking oscillation-induced changes in tunneling current as the membrane resonantly vibrates within tunneling distance. Second, its lateral ($x$/$y$) precision — maintained by dual feedback loops — keeps the tip fixed at the same membrane location throughout, ensuring reproducibility by avoiding spatial variation in local forces.
@@
By setting fixed values for the tunneling current and bias voltage, the system controls the distance between the tip apex and the membrane. When $V_{\rm AC}$ is applied to the backgate, the membrane is driven, and its vibrational amplitude increases the most as the AC signal approaches the membrane’s mechanical resonance frequency. To maintain a constant averaged tunneling current, the STM tip retracts, adjusting its $z$-position via the piezoceramic tube. This $z$-position is then recorded by the control system, providing a direct measure of the membrane’s vibrational amplitude. When the AC signal frequency nears the resonant frequency of the nanomembrane, the increased vibrational amplitude is clearly resolved in the tip’s $z$-position, thereby indicating both the resonance position and the mode amplitude as seen in Fig. 5c.
The STM offers several key advantages for this unconventional application. Unlike optical or capacitive readouts, which can perturb the membrane over large areas and billions of atoms, STM tunneling is extremely localized and equally effective above and below $T_{C}$, making it ideal for studying the superconducting transition. Furthermore, the ultra-high vacuum (UHV) environment required for cryogenic STM operation provides exceptional control over surface conditions. Our STM has annealing capabilities that reduce surface oxidation — an often-overlooked factor in cryogenic nanomechanical setups. NbTiN resists oxidation far better than aluminum [29, 16, 33] due to its stable nitride structure, which limits oxygen diffusion. Unlike aluminum oxides, which are chemically stable and difficult to remove, NbTiN oxidation remains minimal and can be effectively annealed in ultra-high vacuum at 200C, ensuring a cleaner superconducting surface for STM measurements and removing a source of electrostatic patch potentials [12, 13]. The STM tip can directly probe superconducting properties by identifying the superconducting gap and $T_{C}$, offering a complementary diagnostic tool alongside mechanical measurements. In addition to its high $T_{C}$, these characteristics highlight NbTiN’s suitability as a membrane material — its surface compatibility with STM enables robust coupling between the mechanical structure and the tunneling readout. Casimir forces are inherently surface-dominated, as are many noise sources, making STM an ideal probe; here, we demonstrate its dual role in both mechanical detection and direct surface characterization, offering insights inaccessible to conventional methods.
Our measurement protocol is designed to detect changes in the Casimir force across the superconducting transition by tracking shifts in the mechanical resonance frequency of suspended membranes. Rather than measuring absolute Casimir forces — which are subject to large theoretical uncertainties due to unknown superconducting corrections and thermal contributions — we focus on relative changes across $T_{C}$. This differential approach enhances sensitivity to superconductivity-induced modifications of the Casimir interaction. Small-gap (190 nm) and big-gap (1213 nm) configurations are compared to isolate Casimir-specific effects, with the big-gap device serving as a control to subtract out non-Casimir contributions such as thermal expansion, changes in the material, and changes in the tunneling across $T_{C}$. STM’s high spatial precision enables stable, reproducible readout on both devices, providing high-confidence removal of non-Casimir effects.
@@
The experimental setup consists of a loadlock, a preparation chamber, and a measurement chamber. Samples are first loaded into the loadlock and annealed at 200C under a pressure of $1.7\times 10^{-8}$ mbar for two days to remove contaminants. They are then transferred to the preparation and measurement chambers, where experiments are performed at 4.45 K and $1.5\times 10^{-9}$ mbar. As shown in Fig. 3b, the sample holder accommodates chips with both small- and big-gap cavities, allowing side-by-side comparison. AC and DC signals are applied to the backgate beneath the membranes to drive mechanical oscillations and cancel electrostatic forces. A bias voltage on the suspended membrane enables tunneling current detection by the STM tip. Tip placement is guided by an in-situ optical camera (Fig. 3b), enabling coarse positioning near the membrane edge (60 microns away) where vibrational amplitude is minimized. Resonance frequency measurements are then performed at different temperatures across $T_{C}$.
@@
By combining our novel STM readout with microchip-based Casimir cavities, we establish a powerful platform for minimally invasive, high-sensitivity measurements across the superconducting transition. Achieving meaningful results, however, requires careful isolation of all extraneous forces that could obscure the Casimir effect—including those introduced by the STM, electrostatic interactions between the plates, and thermal expansion or unknown material changes at $T_{C}$ during temperature sweeps. In the following section, we outline the methods used to systematically eliminate these contributions, allowing us to isolate signals attributable to the Casimir force.
VI.2 Experimentally Isolating Casimir Forces
Figure 5: Eliminating non-Casimir Effects with STM-Microchip Approach (a) Atomic-scale scan (3×3(3$\times$ 3( 3 × 3 nm2) showing the surface topology of the NbTiN membranes, demonstrating the STM’s ability to achieve sub-atomic spatial resolution (down to 10 pm, smaller than the diameter of an atom) during measurements. (b) STM positioned above the Casimir cavity membrane, requiring the elimination of forces associated with the tip-sample interaction (red, green), electrostatics between the plates (blue), and thermal-elastic effects from temperature sweeps (purple). All forces are canceled in small- and big-gaps device as plotted in the following: (c) Frequency response as a function of backgate voltage to compensate for electrostatic forces between the plates by finding the apex of the parabola. Subplots in the panel were manually spaced for clarity. (d) Frequency response as a function of tip-membrane bias voltage, used to eliminate residual electrostatic forces between the tip and membrane. The shaded areas (yellow) show the range of the frequency shift, which are presented in the insets. (e) Frequency response of membranes over small (top) and big (bottom) gaps as a function of current, used to minimize van der Waals forces between the tip and the membrane. Subplots in the panel were manually spaced for clarity. (f) Final stage: performing large span temperature sweeps to attain elastic contributions from thermal expansion coefficients.

### v2

@@ III Experimental Setup and Measurement Protocol
To isolate superconductivity-induced Casimir shifts from material effects and thermal expansion, we use two chips with distinct gaps: a small-gap cavity (190 nm) where Casimir forces are strong, and a large-gap cavity (1213 nm) where they are negligible (Fig. 3d). Mounting both on the same holder (Fig. 3b) ensures identical environmental conditions, enabling subtraction of material-dependent effects. Although the NbTiN films are co-deposited, we additionally verify that both chips exhibit matched mechanical frequencies and $T_{C}$ (see Supplementary Information A). This differential approach allows us to subtract shared material responses across $T_{C}$ and isolate Casimir forces. All other extraneous forces are actively canceled or suppressed, leaving temperature sweeps around $T_{C}$ to reveal the Casimir contribution.
Measurements Across $T_{C}$
Figure 4: STM-based Casimir measurement and cancellation of non-Casimir forces. (a) Schematic of thermo-elastic, electrostatic, and van der Waals interactions to be removed. (b) STM scan showing tip scanning within a $3\times 3$ nm2 area at a topographic maximum with 10 pm localization. (c–e) Sequential cancellation of plate–plate electrostatics, tip–sample electrostatics, and van der Waals forces. (f) Large- and small-gap devices share $T_{C}\approx 14.2$ K; total resistance including non-superconducting leads (g) Small-gap device shows sharper frequency changes near $T_{C}$. (h) Background-subtracted Casimir pressure gradient agrees in magnitude and sign with plasma–BCS theory (See Methods).
Our platform is designed to detect subtle changes in the Casimir force during slow, controlled sweeps across the superconducting transition temperature $T_{C}$. Resolving effects at the micro-pascal level requires exceptional frequency stability, high force sensitivity, and suppression of environmental noise. As shown in Fig. 4a, isolating the Casimir contribution involves canceling electrostatic interactions between the membrane and backgate ($F_{\mathrm{ES},bg}$), electrostatic and van der Waals forces between the STM tip and membrane ($F_{\mathrm{ES},tip}$, $F_{\mathrm{vdW},tip}$), and elastic changes in the membrane ($F_{\mathrm{elastic}}$). The latter are removed using a co-fabricated large-gap cavity with negligible Casimir interaction.
To probe superconductivity-induced changes, we heat the membranes from 4.45 K to just above $T_{C}\sim 14.2$ K in stabilized steps. An open-flow helium system provides vibration-free cooling for long sweeps, while UHV preserves clean surfaces and minimizes mechanical damping. Each sweep lasts up to 16 hours, with the STM tip locked to within $\sim$10 pm to ensure reproducible readout.
As shown in Fig. 4b, measurements begin with a $3\times 3$ nm2 atomic-scale scan to identify a topographic maximum. The STM tip is then locked to this site with $\sim$10 pm stability, minimizing position-dependent forces and ensuring remarkably reproducible conditions.This keeps the tunneling interaction confined to the same few atoms of the moving membrane throughout the full sweep across $T_{C}$ That atomic-scale control makes STM uniquely suited for precision Casimir measurements, because probe-induced contributions can be quantified and removed cleanly to isolate the Casimir signal.
Figures 4c–e show sequential cancellation of competing forces. Backgate voltages are swept to cancel $F_{\mathrm{ES},bg}$, with no temperature dependence observed. Tip–membrane biases are adjusted to cancel $F_{\mathrm{ES},tip}$, with residual shifts far below the $\sim$0.5 Hz Casimir signal expected from theory. Van der Waals interactions are excluded by verifying that resonance frequency is independent of tunneling current. With parasitic forces minimized, Casimir changes can be isolated. See Methods for more detailed description of cancellation procedures.
To eliminate material-dependent effects, we differentially measure co-fabricated small-gap (190 nm) and large-gap (1213 nm) cavities mounted on the same holder. Both show the same superconducting transition temperature $T_{C}\approx 14.2$ K (Fig. 4f), confirming identical superconducting on top of matching mechanical behavior (see Supplementary Information A). The relative frequency shift $\Delta f$ between the two devices (Fig. 4g) isolates the Casimir contribution and reveals an abrupt change in the Casimir pressure gradient near $T_{C}$.
[…]
Taken together, these measurements establish that the platform operates in the sensitivity regime required to probe superconductivity-dependent Casimir variations across $T_{C}$, while suppressing known competing contributions in situ. The system combines atomic-scale tip localization, high force responsivity, and systematic cancellation of electrostatic, van der Waals, and thermal backgrounds.
The noise-limited frequency resolution is $\langle\delta f\rangle\approx 4.7$ mHz, extracted from Lorentzian fits to the resonance peak using $\langle\delta f\rangle\approx\frac{f_{0}}{2Q\cdot\mathrm{SNR}}$, where $f_{0}$ is the membrane resonance frequency and SNR the signal-to-noise ratio (see Supplementary Information F4 and Ref. [27]). This demonstrated $\langle\delta f\rangle$ is nearly two orders of magnitude below the 0.29 Hz shift predicted by the plasma–BCS model. Simulations yield Casimir responsivities of 1.73 Hz/nN and 0.87 Hz/mPa, placing our device among the most responsive nanoresonator-based Casimir platforms reported to date [28, 29]. In pressure units, this corresponds to sensitivity at the level of tens of micro-Pascals, well below the milli-Pascal scale of leading predictions.
IV Conclusions
This work establishes a new experimental frontier for Casimir measurements, with unprecedented parallelism, sensitivity, and suppression of extraneous forces in cryogenic environments, where less than a handful of Casimir experiments have so far ventured. By combining micro-electromechanical systems (MEMS) cavity engineering with atomic-scale STM readout, we realize ultra-parallel, compliant, high-$Q$ superconducting membranes and differential cancellation protocols that access the micro-Pascal regime across $T_{C}$. This architecture brings Casimir measurements closer to the original parallel-plate limit under superconducting conditions, reducing geometric assumptions and enabling more direct comparisons between theory and experiment in regimes where sphere-plate approximations may be insufficient.
Because superconductivity is a second-order (continuous) phase transition that primarily modifies low-frequency electrodynamics, it has long been unclear whether crossing $T_{C}$ should produce any abrupt or measurable change in the Casimir force [30, 16]. If confirmed as Casimir in origin, such shifts could open a rich field of new physics where superconductivity modifies vacuum fluctuations. The apparent discontinuity, in both direction and magnitude, qualitatively matches with the plasma–BCS model presented in Methods, which uniquely predicts a jump across $T_{C}$. Both the plasma model and BCS theory describe dissipationless electronic responses, but in different senses: the plasma model is dissipationless only by neglecting relaxation in a free-electron gas, while BCS dissipationlessness arises from quantum coherence of the superconducting condensate, which also enforces the Meissner effect. Recent theory and experimental work suggest the plasma model better captures low-frequency transverse electric responses in the normal state [31], while coherence and non-local effects may further modify this behavior [32], highlighting importance of experiments in this uncharted territory of physics.
Looking ahead, the next crucial step is to resolve potential shifts at $T_{C}$ and determine their evolution by systematically varying plate separation. This will require series of nominally identical chips with progressively tuned gaps, extending beyond the two extremes probed here: smallest gaps where Casimir forces dominate and largest gaps where they are negligible. A further step is to separate competing theories using their distinct scaling with parameters such as plate thickness and superconducting $T_{C}$ [33]. Establishing this baseline would move the field beyond long-standing debates and enable credible tests of more speculative directions, from mechanisms in high-$T_{C}$ superconductors [34, 35, 36, 37] to possible gravitational signatures [38, 39, 40, 41, 42]. Any confirmed signal along these directions would be a major advance in our understanding of quantum materials and vacuum energy. Our platform sets a new benchmark for entering this regime and provides a foundation for future discoveries.
V Methods
V.1 On-Chip Casimir Cavity Fabrication
[…]
Figure 5a outlines the fabrication process. A completed 10 × 10 mm2 chip (Fig. 5b) hosts multiple Casimir cavities, each formed from a suspended NbTiN membrane spanning 700 µm over a 190 nm vacuum gap. The scale illustration shows the extreme aspect ratio; the membrane’s central deflection is not visible but is calculated to be only $\sim$153 pm (See Supplementary Information D), set almost entirely by the Casimir force, and would require roughly a $10^{3}$ zoom beyond the gap view to visualize.
Our platform achieves area-to-gap ratios exceeding prior Casimir configurations by orders of magnitude, enabled by monolithic chip-based parallel plates rather than active alignment. This yields nanometer-scale gaps over hundreds of micrometers with robustness, scalability, and reproducibility. Stability, compliance, and smoothness are engineered directly into the structure using high-stress superconductors, sacrificial amorphous silicon, and cryogenic plasma etching.
Layered chips are produced by sequential deposition: thermal oxide for insulation, a bottom NbTiN electrode, an ICP-CVD amorphous silicon sacrificial layer, and a top NbTiN membrane, followed by annealing and dicing. The flat, tensile trilayer geometry avoids thermally induced out-of-plane distortions seen in drum-type membranes, simplifying temperature sweeps and analysis.
Casimir cavities are defined by etching release holes in the top NbTiN via electron beam lithography and fluorine plasma, then removing the a-Si with a cryogenic SF6 plasma undercut at –120 C. The etch uses ICP power only (no RF bias) to protect the superconducting membrane; low temperature slows NbTiN etching, preserves extreme aspect ratios, and improves surface smoothness. Devices are then warmed slowly and adiabatically from cryogenic temperatures to room temperature under vacuum, an essential step to prevent collapse, before being transferred to the STM lab. Even slight temperature gradients across a strained membrane can cause deformations that can lead to runaway collapse. This was a key ingredient for reliably fabricating smaller gap devices.
Two key innovations enable this platform: (1) ICP-CVD deposition of low-hydrogen a-Si with high uniformity, ensuring predictable gaps and avoiding blistering; and (2) cryogenic SF6 etching with $\sim$1000:1 selectivity to NbTiN, releasing ultra-compliant membranes without damage. Together, these methods allow large-span metallic resonators to be suspended at Casimir-dominated gaps without active alignment, eliminating drift and enabling quantum-limited parallelism for Casimir studies at any temperature.
V.2 STM measurement protocol
The STM offers several key advantages for this application. Unlike optical or capacitive readouts, which perturb the membrane over large areas and billions of atoms, STM tunneling is highly localized and equally effective above and below $T_{C}$, making it suitable for studying the superconducting transition. The ultra-high vacuum (UHV) environment required for cryogenic STM operation provides strong control over surface conditions. Our STM setup includes annealing capabilities that reduce surface oxidation, an often-overlooked factor in cryogenic nanomechanical setups. NbTiN resists oxidation more effectively than aluminum [18, 15, 23] due to its stable nitride structure, which limits oxygen diffusion. Unlike aluminum oxides, which are chemically stable and difficult to remove, NbTiN oxidation is minimal and can be annealed in UHV at 200 C, reducing electrostatic patch potentials [12, 13] and improving tunneling current signal. The STM tip can also probe superconducting properties by identifying the superconducting gap and $T_{C}$, providing a diagnostic tool alongside mechanical measurements (See Supplementary Information E). In addition to its high $T_{C}$, NbTiN’s surface compatibility with STM enables coupling between the membrane and tunneling readout. Casimir forces are surface-dominated, as are many noise sources, making STM well-suited for both mechanical detection and surface characterization.
Our protocol detects changes in the Casimir force across the superconducting transition by tracking shifts in the mechanical resonance frequency of suspended membranes. Instead of measuring absolute Casimir forces, which carry large theoretical uncertainties from unknown superconducting corrections and thermal effects, we measure relative changes across $T_{C}$. This differential approach is more sensitive to superconductivity-induced modifications of the Casimir interaction. Small-gap (190 nm) and large-gap (1213 nm) devices are compared; the large-gap device serves as a control to subtract non-Casimir contributions such as thermal expansion, material changes, and tunneling variations across $T_{C}$. STM’s high spatial precision enables stable, reproducible readout on both devices, ensuring confident removal of non-Casimir effects.
@@ V.2 STM measurement protocol
As described in the main text, the STM provides high-sensitivity, tunneling current readout of membrane resonance across the superconducting transition. Here, we detail the protocol used to achieve the stability, localization, and noise suppression required for the results in Fig. 4. We use a scanning tunneling microscope (STM) to measure membrane dynamics with atomic-scale precision. The tunneling current depends exponentially on tip–sample separation, with 0.1 nm changes producing an order-of-magnitude variation. This extreme sensitivity allows us to track the mechanical resonance of suspended superconducting NbTiN membranes, extending STM from atomic-scale imaging to mesoscopic devices.
@@
Two aspects of the STM’s precision are essential. First, its high vertical ($z$) resolution detects membrane resonance by tracking oscillation-induced changes in tunneling current within tunneling distance. Second, its lateral (xx/$y$) precision, maintained by dual feedback loops, keeps the tip fixed at the same membrane location, avoiding spatial variation in local forces.
@@
By setting fixed values for the tunneling current and bias voltage, the system controls the distance between the tip apex and the membrane. When $V_{\rm AC}$ is applied to the backgate, the membrane is driven, and its vibrational amplitude peaks as the AC signal approaches the membrane’s mechanical resonance frequency. To maintain a constant averaged tunneling current, the STM tip retracts, adjusting its $z$-position via the piezoceramic tube. This $z$-position is recorded by the control system as a direct measure of the membrane’s vibrational amplitude. When the AC signal frequency nears the resonant frequency of the nanomembrane, the increased vibrational amplitude is clearly resolved in the tip’s $z$-position, indicating both the resonance position and the mode amplitude as seen in Fig. 4e.
@@
The experimental setup consists of a loadlock, a preparation chamber, and a measurement chamber. Samples are first loaded into the loadlock and annealed at 200C under a pressure of $1.7\times 10^{-8}$ mbar for two days to remove contaminants. They are then transferred to the preparation and measurement chambers, where experiments are performed at 4.45 K and $1.5\times 10^{-10}$ mbar. As shown in Fig. 3b, the sample holder accommodates chips with both small- and big-gap cavities, allowing side-by-side comparison. AC and DC signals are applied to the backgate beneath the membranes (via wire bonding) to drive mechanical oscillations and cancel electrostatic forces. A bias voltage on the suspended membrane enables tunneling current detection by the STM tip. Tip placement is guided by an in-situ optical camera (Fig. 3c), enabling coarse positioning near the membrane edge (60 microns away) where vibrational amplitude is minimized. Resonance frequency measurements are then performed at different temperatures across $T_{C}$.
@@
By combining STM readout with microchip-based Casimir cavities, we create a platform for minimally invasive, high-sensitivity measurements across the superconducting transition. To obtain meaningful results, it is essential to remove all extraneous forces that could mask the Casimir effect, including those from the STM, electrostatic interactions between the plates, and changes from thermal expansion or material properties at $T_{C}$ during temperature sweeps. While the main text describes these cancellation strategies conceptually, the following subsections detail their exact implementation in our experiments.
V.3 Experimentally isolating Casimir forces